OpenStudy (anonymous):

solve the quadratic 2cos^2θ - cosθ =1

OpenStudy (anonymous):

you can simply: \[2\cos^2\theta-\cos \theta-1=0\] Now treat this as a quadratic, simply let cos theta be represented by "a": \[2a^2-a-1=0\] Then factor: \[(2a+1)(a-1)=0\] Next set each factor to zero: \[2a+1=0\] and \[a-1=0\] Which means that : \[a=-1/2\] and a=1 Now simply replace "a" with cos: cos(theta)=-1/2 and cos(theta)=1 Now just solve for the angles

OpenStudy (anonymous):

thankyou!